\begin{thebibliography}{99}

  %\cite{Burkardt:2003yg}
\bibitem{Burkardt:2003yg}
  M.~Burkardt,
  %``Quark correlations and single spin asymmetries,''
  Phys.\ Rev.\ D {\bf 69}, 057501 (2004)
  [hep-ph/0311013].
  %%CITATION = HEP-PH/0311013;%%
  %79 citations counted in INSPIRE as of 25 Feb 2015

  %\cite{Sivers:1989cc}
\bibitem{Sivers:1989cc}
  D.~W.~Sivers,
  %``Single Spin Production Asymmetries from the Hard Scattering of Point-Like Constituents,''
  Phys.\ Rev.\ D {\bf 41}, 83 (1990).
  %%CITATION = PHRVA,D41,83;%%
  %818 citations counted in INSPIRE as of 25 Feb 2015


%\cite{Lorce:2015lna}
\bibitem{Lorce:2015lna}
  C.~Lorc\'{e},
  %``The light-front gauge-invariant energy-momentum tensor,''
  arXiv:1502.06656 [hep-ph].
  %%CITATION = ARXIV:1502.06656;%%


  %\cite{Goeke:2006ef}
\bibitem{Goeke:2006ef}
  K.~Goeke, S.~Meissner, A.~Metz and M.~Schlegel,
  %``Checking the Burkardt sum rule for the Sivers function by model calculations,''
  Phys.\ Lett.\ B {\bf 637}, 241 (2006)
  [hep-ph/0601133].
  %%CITATION = HEP-PH/0601133;%%
  %23 citations counted in INSPIRE as of 25 Feb 2015

%\cite{Courtoy:2008vi}
\bibitem{Courtoy:2008vi}
  A.~Courtoy, F.~Fratini, S.~Scopetta and V.~Vento,
  %``A Quark model analysis of the Sivers function,''
  Phys.\ Rev.\ D {\bf 78}, 034002 (2008)
  [arXiv:0801.4347 [hep-ph]].
  %%CITATION = ARXIV:0801.4347;%%
  %44 citations counted in INSPIRE as of 25 Feb 2015

%\cite{Courtoy:2008dn}
\bibitem{Courtoy:2008dn}
  A.~Courtoy, S.~Scopetta and V.~Vento,
  %``Model calculations of the Sivers function satisfying the Burkardt Sum Rule,''
  Phys.\ Rev.\ D {\bf 79}, 074001 (2009)
  [arXiv:0811.1191 [hep-ph]].
  %%CITATION = ARXIV:0811.1191;%%
  %37 citations counted in INSPIRE as of 25 Feb 2015

  %\cite{Pasquini:2010af}
\bibitem{Pasquini:2010af}
  B.~Pasquini and F.~Yuan,
  %``Sivers and Boer-Mulders functions in Light-Cone Quark Models,''
  Phys.\ Rev.\ D {\bf 81}, 114013 (2010)
  [arXiv:1001.5398 [hep-ph]].
  %%CITATION = ARXIV:1001.5398;%%
  %45 citations counted in INSPIRE as of 25 Feb 2015

%\cite{Pasquini:2011tk}
\bibitem{Pasquini:2011tk}
  B.~Pasquini and P.~Schweitzer,
  %``Naive time-reversal odd phenomena in semi-inclusive deep-inelastic scattering from light-cone constituent quark models,''
  Phys.\ Rev.\ D {\bf 83}, 114044 (2011)
  [arXiv:1103.5977 [hep-ph]].
  %%CITATION = ARXIV:1103.5977;%%
  %27 citations counted in INSPIRE as of 25 Feb 2015



  %\cite{Boer:2003cm}
\bibitem{Boer:2003cm}
  D.~Boer, P.~J.~Mulders and F.~Pijlman,
  %``Universality of T odd effects in single spin and azimuthal asymmetries,''
  Nucl.\ Phys.\ B {\bf 667}, 201 (2003)
  [hep-ph/0303034].
  %%CITATION = HEP-PH/0303034;%%
  %395 citations counted in INSPIRE as of 25 Feb 2015

       %\cite{Efremov:1981sh}
\bibitem{Efremov:1981sh}
  A.~V.~Efremov and O.~V.~Teryaev,
  %``On Spin Effects In Quantum Chromodynamics,''
  Sov.\ J.\ Nucl.\ Phys.\  {\bf 36}, 140 (1982)
  [Yad.\ Fiz.\  {\bf 36}, 242 (1982)];
  %%CITATION = SJNCA,36,140;%%
%\cite{Efremov:1984ip}
%\bibitem{Efremov:1984ip}
  %``QCD Asymmetry And Polarized Hadron Structure Functions,''
  Phys.\ Lett.\ B {\bf 150}, 383 (1985).
  %%CITATION = PHLTA,B150,383;%%

%\cite{Qiu:1991pp}
\bibitem{Qiu:1991pp}
  J.-w.~Qiu and G.~F.~Sterman,
  %``Single transverse spin asymmetries,''
  Phys.\ Rev.\ Lett.\  {\bf 67}, 2264 (1991);
  %CITATION = PRLTA,67,2264;%%

  %\cite{Ji:1992eu}
\bibitem{Ji:1992eu}
  X.~-D.~Ji,
  %``Gluon correlations in the transversely polarized nucleon,''
  Phys.\ Lett.\ B {\bf 289}, 137 (1992).
  %%CITATION = PHLTA,B289,137;%%
  %76 citations counted in INSPIRE as of 05 Mar 2013


  %\cite{Kang:2008qh}
\bibitem{Kang:2008qh}
  Z.~-B.~Kang and J.~-W.~Qiu,
  %``Single transverse-spin asymmetry for D-meson production in semi-inclusive deep inelastic scattering,''
  Phys.\ Rev.\ D {\bf 78}, 034005 (2008)
  [arXiv:0806.1970 [hep-ph]].
  %%CITATION = ARXIV:0806.1970;%%
  %40 citations counted in INSPIRE as of 05 Mar 2013

    %\cite{Beppu:2010qn}
\bibitem{Beppu:2010qn}
  H.~Beppu, Y.~Koike, K.~Tanaka and S.~Yoshida,
  %``Contribution of Twist-3 Multi-Gluon Correlation Functions to Single Spin Asymmetry in Semi-Inclusive Deep Inelastic Scattering,''
  Phys.\ Rev.\ D {\bf 82}, 054005 (2010)
  [arXiv:1007.2034 [hep-ph]].
  %%CITATION = ARXIV:1007.2034;%%
  %25 citations counted in INSPIRE as of 05 Mar 2013



%\cite{Kang:2008ey}
\bibitem{Kang:2008ey}
  Z.~-B.~Kang and J.~-W.~Qiu,
  %``Evolution of twist-3 multi-parton correlation functions relevant to single transverse-spin asymmetry,''
  Phys.\ Rev.\ D {\bf 79}, 016003 (2009).
  %%CITATION = ARXIV:0811.3101;%%



  %\cite{Zhou:2008mz}
\bibitem{Zhou:2008mz}
  J.~Zhou, F.~Yuan and Z.~-T.~Liang,
  %``QCD Evolution of the Transverse Momentum Dependent Correlations,''
  Phys.\ Rev.\ D {\bf 79}, 114022 (2009).
  %%CITATION = ARXIV:0812.4484;%%



  %\cite{Vogelsang:2009pj}
\bibitem{Vogelsang:2009pj}
  W.~Vogelsang and F.~Yuan,
  %``Next-to-leading Order Calculation of the Single Transverse Spin Asymmetry in the Drell-Yan Process,''
  Phys.\ Rev.\ D {\bf 79}, 094010 (2009).
  %%CITATION = ARXIV:0904.0410;%%


  %\cite{Braun:2009mi}
\bibitem{Braun:2009mi}
  V.~M.~Braun, A.~N.~Manashov and B.~Pirnay,
  %``On the scale dependence of twist-three contributions to single spin asymmetries,''
  Phys.\ Rev.\ D {\bf 80}, 114002 (2009).
  %%CITATION = ARXIV:0909.3410;%%

  %\cite{Ma:2011nd}
\bibitem{Ma:2011nd}
  J.~P.~Ma and H.~Z.~Sang,
  %``Soft-Gluon-Pole Contribution in Single Transverse-Spin Asymmetries of Drell-Yan Processes,''
  JHEP {\bf 1104}, 062 (2011)
  [arXiv:1102.2679 [hep-ph]].
  %%CITATION = ARXIV:1102.2679;%%
  %7 citations counted in INSPIRE as of 05 Mar 2013

  %\cite{Schafer:2012ra}
\bibitem{Schafer:2012ra}
  A.~Schafer and J.~Zhou,
  %``A Note on the scale evolution of the ETQS function $T_F(x,x)$,''
  Phys.\ Rev.\ D {\bf 85}, 117501 (2012)
  [arXiv:1203.5293 [hep-ph]].
  %%CITATION = ARXIV:1203.5293;%%
  %18 citations counted in INSPIRE as of 25 Feb 2015

    %\cite{Ma:2012xn}
\bibitem{Ma:2012xn}
  J.~P.~Ma and Q.~Wang,
  %``Scale Dependence of Twist-3 Quark-Gluon Operators for Single Spin Asymmetries,''
  Phys.\ Lett.\ B {\bf 715}, 157 (2012)
  [arXiv:1205.0611 [hep-ph]].
  %%CITATION = ARXIV:1205.0611;%%
  %8 citations counted in INSPIRE as of 05 Mar 2013

%\cite{Kang:2012em}
\bibitem{Kang:2012em}
  Z.~B.~Kang and J.~W.~Qiu,
  %``QCD evolution of naive-time-reversal-odd parton distribution functions,''
  Phys.\ Lett.\ B {\bf 713}, 273 (2012)
  [arXiv:1205.1019 [hep-ph]].
  %%CITATION = ARXIV:1205.1019;%%
  %22 citations counted in INSPIRE as of 10 juil. 2015

%\cite{Sun:2013hua}
\bibitem{Sun:2013hua}
  P.~Sun and F.~Yuan,
  %``Transverse momentum dependent evolution: Matching semi-inclusive deep inelastic scattering processes to Drell-Yan and W/Z boson production,''
  Phys.\ Rev.\ D {\bf 88}, no. 11, 114012 (2013)
  [arXiv:1308.5003 [hep-ph]].
  %%CITATION = ARXIV:1308.5003;%%
  %31 citations counted in INSPIRE as of 25 Feb 2015

%\cite{Schafer:2013wca}
\bibitem{Schafer:2013wca}
  A.~Schafer and J.~Zhou,
  %``Transverse single spin asymmetry in hadronic $\eta_{c,b}$ production,''
  Phys.\ Rev.\ D {\bf 88}, no. 1, 014008 (2013)
  [arXiv:1302.4600 [hep-ph]].
  %%CITATION = ARXIV:1302.4600;%%
  %3 citations counted in INSPIRE as of 25 Feb 2015

%\cite{Schafer:2013opa}
\bibitem{Schafer:2013opa}
  A.~Schafer and J.~Zhou,
  %``A note on the scale evolution of tri-gluon correlations,''
  arXiv:1308.4961 [hep-ph].
  %%CITATION = ARXIV:1308.4961;%%
  %5 citations counted in INSPIRE as of 25 Feb 2015

%\cite{Dai:2014ala}
\bibitem{Dai:2014ala}
  L.~Y.~Dai, Z.~B.~Kang, A.~Prokudin and I.~Vitev,
  %``Next-to-leading order transverse momentum-weighted Sivers asymmetry in semi-inclusive deep inelastic scattering: the role of the three-gluon correlator,''
  arXiv:1409.5851 [hep-ph].
  %%CITATION = ARXIV:1409.5851;%%

      %\cite{Ratcliffe:2014nla}
\bibitem{Ratcliffe:2014nla}
  P.~G.~Ratcliffe and O.~V.~Teryaev,
  %``On the Evolution of Sum Rules for T-Odd Distribution and Fragmentation Functions,''
  arXiv:1406.1604 [hep-ph].
  %%CITATION = ARXIV:1406.1604;%%
  %1 citations counted in INSPIRE as of 16 Feb 2015

  %\cite{Zhou:2013gsa}
\bibitem{Zhou:2013gsa}
  J.~Zhou,
  %``Transverse single spin asymmetries at small x and the anomalous magnetic moment,''
  Phys.\ Rev.\ D {\bf 89}, no. 7, 074050 (2014)
  [arXiv:1308.5912 [hep-ph]].
  %%CITATION = ARXIV:1308.5912;%%
  %3 citations counted in INSPIRE as of 25 Feb 2015

\end{thebibliography}

\documentclass[prd,aps,preprint,nofootinbib,superscriptaddress]{revtex4}
\usepackage{epsfig}
\usepackage{amsmath}
\input{epsf}

\begin{document}


\vspace*{2cm}
\title{A note on the scale dependence of the Burkardt sum rule}

\author{Jian Zhou}
\affiliation{\normalsize\it Nikhef and Department of Physics and Astronomy, VU University Amsterdam,  De Boelelaan 1081, NL-1081 HV Amsterdam, The Netherlands}

\begin{abstract}
In this short note, we argue that the Burkardt sum rule for the Sivers functions
can be used to check the consistency of evolution equations of three-parton correlators.
\end{abstract}

\maketitle

Sum rules are often very useful for providing model independent constraints for non-perturbative objects.
Among them, the Burkardt sum rule~\cite{Burkardt:2003yg} is of particular interest in studying transverse single spin asymmetry(SSA) phenomenology.
It states that the net transverse momentum carried by partons inside a transversely polarized nucleon
vanishes when summing over all parton flavors,
\begin{eqnarray}
\sum_{a=q, \bar q, g} \langle k_\perp^a \rangle =0
\end{eqnarray}
Expressed in terms of the Sivers functions~\cite{Sivers:1989cc} it takes form,
\begin{eqnarray}
\sum_{a=q, \bar q, g} \int_0^1 dx \int d^2  k_\perp \  k_\perp^2 f_{1T}^{\perp a}(x,k_\perp^2)=0
\end{eqnarray}
The Burkardt sum rule was first derived in Ref.~\cite{Burkardt:2003yg},
 and previously has been proved in an alternative way by Lorce~\cite{Lorce:2015lna}.
 It has been checked in various
 model calculations~\cite{Goeke:2006ef,Courtoy:2008vi,Courtoy:2008dn,Pasquini:2010af,Pasquini:2011tk}.

Using the well known tree level relations between the $ k_\perp^2$ moment of the Sivers functions
and the corresponding collinear twist-3 correlations~\cite{Boer:2003cm}, the sum rule can be  re-expressed as,
\begin{eqnarray}
\int_0^1 dx T_G^{(+)}(x,x)+\sum_{a=q, \bar q} \int_0^1 dx T_F^a(x,x)=0
\end{eqnarray}
where $T_F$ is the Qiu-Sterman(QS) function~\cite{Efremov:1981sh,Qiu:1991pp},
and $T_G^{(+)}$ is the C-even tri-gluon correlation~\cite{Ji:1992eu,Kang:2008qh,Beppu:2010qn}.
The scale dependence of the sum rule thus can be investigated by using the scale evolution equations of the relevant twist-3
 correlations $T_F$ and $T_G^{(+)}$ which already existed in the  literatures
~\cite{Kang:2008ey,Zhou:2008mz,Vogelsang:2009pj,Braun:2009mi,Ma:2011nd,Schafer:2012ra,Ma:2012xn,Kang:2012em,Sun:2013hua,Schafer:2013wca,Schafer:2013opa,Dai:2014ala}.
 Note that this subject has also been studied from a different aspect of view in Ref.~\cite{Ratcliffe:2014nla}.

We start from the scale evolution equation for quark QS function
$T_F^q(x,x)$~\cite{Kang:2008ey,Zhou:2008mz,Vogelsang:2009pj,Braun:2009mi,Ma:2011nd,Schafer:2012ra,Ma:2012xn,Kang:2012em,Sun:2013hua},
\begin{eqnarray}
&&\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!
\frac{\partial T_F^q(\xi,\xi,\mu^2)}{\partial {\ln\mu^2}}|_{q, \bar q\rightarrow q}
=\frac{\alpha_s}{2\pi} \int_{\xi} \frac{dx}{x} \left [
C_F \left \{ \frac{1+z^2}{(1-z)_+}+ \frac{3}{2} \delta(1-z) \right
\} T_F^q(x,x) \right .\
\nonumber \\
&& + \left .\ \!\!\! \frac{C_A}{2} \left \{
\frac{1+z}{1-z } T_F^q(\xi,x) -\frac{1+z^2}{1-z}  T_F^q(x,x)-2\delta(1-z)T_F^q(x,x) \right \} \right .\
\nonumber \\
&&- \left .\ \!\!\!
\frac{N_c}{2} \tilde T_F^q(\xi,x)+\frac{1}{2N_c} (1-2z) T_F^q(\xi, \xi-x)
-\frac{1}{2N_c}\tilde T_F^q(\xi,\xi-x)
\right ]
\end{eqnarray}
where $z=\xi/x$. The last two terms in the above formula contain anti-quark contribution.
We proceed by carrying out the integration over $\xi$ on both sides of the above equation,
\begin{eqnarray}
\frac{\partial T_F^q(\mu^2)}{\partial {\ln\mu^2}}|_{q, \bar q \rightarrow q}
&=& \!\!
\frac{\alpha_s}{2\pi}   \frac{C_A}{2} \left \{
\int dx  dz \frac{1+z}{1-z } T_F^q(zx,x) -T_F^q(\mu^2)\int dz \frac{1+z^2}{1-z}
\right .\
\nonumber \\ &&
\left .\ \ \ \ \ \ \ \ \ \ \ \
-\int dx  dz \tilde T_F^q(zx,x)
-2T_F^q(\mu^2) \right \}
\nonumber \\ &+& \!\!
\frac{\alpha_s}{2\pi}\frac{1}{2N_c} \int dx  dz \left \{ (1-2z) T_F^q(zx,(z-1)x)
- \tilde T_F^q(zx,(z-1)x) \right \}
\end{eqnarray}
where we introduce the short hand notation $T_F^q(\mu^2)=\int d\xi \ T_F^q(\xi,\xi, \mu^2)$.
The same analysis applies to the scale evolution of anti-quark QS function $T_F^{\bar q}(\mu^2)$.
Note that our convention for $T_F^{\bar q}$ differs from the one used in Ref.~\cite{Kang:2008ey} by a minus sign,
 such that the relation between the $k_\perp^2$ moment of the anti-quark Sivers function and $T_F^{\bar q}$ is the same as the
 relation for the quark ones in a same process(for example, SIDIS process).
Combining quark and anti-quark contributions, one obtains,
\begin{eqnarray}
\frac{\partial \left [ T_F^q(\mu^2)+ T_F^{\bar q}(\mu^2)\right ]}{\partial {\ln\mu^2}}|_{q,\bar q \rightarrow q,\bar q}
&=&
\frac{\alpha_s}{2\pi}   \frac{C_A}{2} \left \{
\int dx  dz \frac{1+z}{1-z } \left [ T_F^q(zx,x)+T_F^{\bar q}(zx,x)] \right ]
\right .\
\nonumber \\ &&    \left .\
 -\int dz \frac{1+z^2}{1-z} \left [ T_F^q(\mu^2)+T_F^{\bar q}(\mu^2) \right ]
 \right .\
 \nonumber \\ & & \left .\ \!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\! -
\int dx  dz \left [\tilde T_F^q(zx,x)+\tilde T_F^{\bar q}(zx,x) \right ]-
 2\left [ T_F^q(\mu^2)+T_F^{\bar q}(\mu^2) \right ] \right \}
\label{qq}
\end{eqnarray}
To arrive the above result, we have made use of the symmetry properties $T_F^{q,\bar q}(x_1,x_2)=T_F^{q,\bar q}(x_2,x_1)$,
 $\tilde T_F^{q,\bar q}(x_1,x_2)=-\tilde T_F^{q,\bar q}(x_2,x_1)$ and the relations $T_F^q(x_1,-x_2)=T_F^{\bar q}(-x_1,x_2)$,
 $\tilde T_F^q(x_1,-x_2)=\tilde T_F^{\bar q} (-x_1,x_2)$~\cite{Kang:2008ey,Braun:2009mi}.

The scale evolution of the tri-gluon correlation $T_G^{(+)}$ also receives the contribution
from quark and anti-quark~\cite{Braun:2009mi,Schafer:2013wca}
\begin{eqnarray}
\frac{\partial T_G^{(+)}(\xi,\xi,\mu^2)}{\partial {\ln\mu^2}}|_{q, \bar q\rightarrow g}
&=&\frac{\alpha_s}{2\pi} \sum_{q ,\bar q} \int_{\xi} \frac{dx}{x} \frac{C_A}{2} \left \{
\frac{1+(1-z)^2}{z } \left [ T_F^q(x,x) +T_F^{\bar q}(x,x)\right ] \right .\
\nonumber \\ && \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \  \left .\
-\frac{2-z}{z}  \left [ T_F^q(x,x-\xi)+T_F^{\bar q}(x,x-\xi) \right ] \right .\
\nonumber \\ &&  \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left .\
+\left [ \tilde T_F^q(x-\xi,x)+\tilde T_F^{\bar q}(x-\xi,x) \right ]
\right \}
\end{eqnarray}
Carrying out the integration over $\xi$ on both sides of the equation and changing the
integration variable $z \rightarrow 1-y$,
\begin{eqnarray}
\frac{\partial T_G^{(+)}(\mu^2)}{\partial {\ln\mu^2}}|_{q, \bar q\rightarrow g}
&=&\frac{\alpha_s}{2\pi}\sum_{q ,\bar q} \frac{C_A}{2}
\left \{ \left [ T_F^q(\mu^2) +T_F^{\bar q}(\mu^2)\right ]
\int dy \frac{1+y^2}{1-y }  \right .\
\nonumber \\ && \ \ \ \ \ \ \ \ \ \
\left .\
+ \int dx \int dy \left (
-\frac{1+y}{1-y}  \right )  \left [ T_F^q(x,yx)+T_F^{\bar q}(x,yx) \right ] \right .\
\nonumber \\ && \ \ \ \ \ \ \ \ \ \
\left .\
+ \int dx \int dy
\left [ \tilde T_F^q(yx,x)+\tilde T_F^{\bar q}(yx,x) \right ]
\right \}
\label{qg}
\end{eqnarray}
Adding up Eq.~\ref{qq} and Eq.~\ref{qg}, we obtain,
\begin{eqnarray}
\frac{\partial \sum_{q ,\bar q} \left [ T_F^q(\mu^2)+ T_F^{\bar q}(\mu^2)+T_G^{(+)}(\mu^2) \right ] }{\partial {\ln\mu^2}}|_{q, \bar q \rightarrow q, \bar q,g}
=- \frac{\alpha_s}{2\pi} \sum_{q ,\bar q} C_A \left [ T_F^q(\mu^2)+ T_F^{\bar q}(\mu^2) \right ]
\label{qqg}
\end{eqnarray}

We now consider the contribution from the tri-gluon correlation
$T_G^{(+)}(x,x)$ to the scale evolution of the  quark  and anti-quark QS functions.
It is worthy to mention that
the O type tri-gluon correlation(related to the function $T_G^{(-)}(x,x)$)  contributes the same in magnitude
and opposite in sign to both the quark and anti-quark Sivers functions.
This is not  surprising  because the O type tri-gluon relation is related to the
$k_\perp$ moment of the spin dependent odderon which measures the difference between particle and anti-particle scattering~\cite{Zhou:2013gsa}.
Therefore, to study the scale dependence of the sum rule, we only need to consider $T_G^{(+)}(x,x)$
contribution which has been given in Refs.~\cite{Kang:2008ey,Ma:2012xn,Dai:2014ala},
\begin{eqnarray}
\frac{\partial T_F^q(\xi,\xi,\mu^2) }{\partial {\rm ln} \mu_F^2} |_{g \rightarrow q}
   &=& \frac{\alpha_s}{2\pi}  \int_\xi \frac{d x}{x} \frac{1}{2} \left[ z^2+(1-z)^2 \right ]
T_G^{(+)}(x,x)
\end{eqnarray}
which leads to
\begin{eqnarray}
\frac{\partial T_F^q(\mu^2) }{\partial {\rm ln} \mu_F^2} |_{g \rightarrow q}
   &=& \frac{\alpha_s}{2\pi} \frac{1}{3}
T_G^{(+)}(\mu^2)
\label{gq}
\end{eqnarray}
The same equation applies to the anti-quark case.

We move on to discuss the contribution from the tri-gluon correlation $T_G^{(+)}$ to the  scale evolution of itself.
Three different evolution equations have been derived by three groups, respectively~\cite{Kang:2008ey,Braun:2009mi,Schafer:2013opa}.
The possible source of the discrepancy between the results in Refs.~\cite{Kang:2008ey,Schafer:2013opa} is that the different parametrization for
the tri-gluon correlations have been used~\cite{Kang:2008qh,Beppu:2010qn}. We are not able to localize the source of the discrepancy between the results~\cite{Braun:2009mi,Schafer:2013opa},
as the authors of paper~\cite{Braun:2009mi} derived the evolution equations in a very different formalism.
Here, we use the one obtained in Ref.~\cite{Schafer:2013opa},
\begin{eqnarray}
\frac{\partial \left [ \frac{N(\xi,\xi)- N(\xi,0)}{\xi} \right ]}{\partial {\rm ln} \mu_F^2} |_{g \rightarrow g}
   &=&
\frac{\alpha_s}{2\pi}C_A  \int_\xi^1 \frac{d x}{x^2}
\left \{
 \frac{(z^2-z+1)^2}{z(1-z)_+} \left [ N(x,x)
 -  N(x,0)\right ] \right .\
\nonumber \\ &&
 + \frac{1+z^2}{2z(1-z)_+} N(\xi,x)
  - \frac{1+(1-z)^2}{2z(1-z)_+} N(x,x-\xi)
    \nonumber \\ && \left .\ \!\!\!
 -\frac{z^2+(1-z)^2}{2z(1-z)_+}N(\xi,\xi-x)
 - \delta(1-z)\left [N(x,x)-N(x,0)\right ] \right \}
 \nonumber \\ && + \frac{\alpha_s}{2\pi}
 \left ( C_A \frac{11}{6}-\frac{n_f}{3} \right ) \left [N(\xi,\xi)-N(\xi,0)\right ]
 \label{gequation}
\end{eqnarray}
where the combination of the N type tri-gluon correlation defined in Ref.~\cite{Beppu:2010qn} can be related to
$T_G^{(+)}$ through the following identity,
\begin{eqnarray}
 \frac{x}{2 \pi} T_G^{(+)}(x,x)&=&-4M_N \left [ N(x,x)- N(x,0) \right ]
\end{eqnarray}
Carrying out the integration over $\xi$ on the both sides of Eq.~\ref{gequation}, we obtain,
\begin{eqnarray}
\frac{T_G^{(+)}(\mu^2)}{\partial {\rm ln} \mu_F^2} |_{g \rightarrow g}
   =
-\frac{\alpha_s}{2\pi}C_A T_G^{(+)}(\mu^2)
-\frac{\alpha_s}{2\pi} \frac{n_f}{3}  T_G^{(+)}(\mu^2)
 \label{gg}
\end{eqnarray}
To arrive the above compact expression, we have used the same mathematic trick, i.e. changing integration variable
$z\rightarrow 1-y$, as well as the symmetry properties for the N type tri-gluon correlation:
$ N(x_1,x_2) = N(x_2, x_1) , \ \ N(x_1,x_2) = -N(-x_1,-x_2)$~\cite{Beppu:2010qn}.

Collecting the intermediate results presented in Eq.~\ref{qqg}, Eq.~\ref{gq}, and Eq.~\ref{gg}, one obtains,
\begin{eqnarray}
\frac{\partial \sum_{q,\bar q,g} \left [ T_F^q(\mu^2)+ T_F^{\bar q}(\mu^2)+T_G^{(+)}(\mu^2) \right ] }{\partial {\ln\mu^2}}
=- \frac{\alpha_s}{2\pi}  C_A \sum_{q,\bar q,g} \left [ T_F^q(\mu^2)+ T_F^{\bar q}(\mu^2) +T_G^{(+)}(\mu^2)\right ]
\end{eqnarray}
It is worthy to mention that all terms on the right side of the equation entirely
come from the so-called boundary terms first discovered by Braun, Manashov and Pirnay~\cite{Braun:2009mi}.
We reexpress the above formula as,
\begin{eqnarray}
\frac{\partial  \left [ \sum_{a=q, \bar q, g} \langle k_\perp^a \rangle (\mu^2) \right ] }{\partial {\ln\mu^2}}
 =- \frac{\alpha_s}{2\pi} C_A \sum_{a=q, \bar q, g} \langle k_\perp^a \rangle(\mu^2)
\end{eqnarray}
which can be readily solved,
\begin{eqnarray}
\sum_{a=q, \bar q, g} \langle k_\perp^a \rangle (Q^2)
= e^{-\frac{C_A}{2\pi}\int_{Q_0^2}^{Q^2} \alpha_s(\mu^2) \frac{d \mu^2}{ \mu^2 }}
\sum_{a=q, \bar q, g} \langle k_\perp^a \rangle (Q^2_0)
\label{final}
\end{eqnarray}
This is the main result of our short note.
It is easy to see that the Burkardt sum rule holds at any scale provided that it is satisfied  at one
arbitrary scale. If the Burkardt sum rule were violated at a low initial  scale for some unknown reason, judging from the
above formula, it will be asymptotically  satisfied at higher scale.

To summarize: We found that the Burkardt sum rule is stable under QCD
scale evolution if one uses the  tri-gluon evolution equation obtained in Ref.~\cite{Schafer:2013opa}.
However, without making any model dependent assumption for the function $T_G^{(+)}$, we are not able to show that the Burkardt sum rule is  preserved under  QCD scale evolution when
the different tri-gluon evolution equations are used~\cite{Kang:2008ey,Braun:2009mi}.
These findings might shed new light on the controversy  about the tri-gluon scale evolution.

\

\noindent
{\bf Acknowledgments:}
I am grateful to A. Metz for initiating this work and the critical reading of the manuscript.
I also would like to thank D. Boer and P. Mulders for helpful discussions and valuable  comments.
This research has been supported by the EU "Ideas" program QWORK (contract 320389).

\end {document}